OpenStudy (anonymous):

Convert to proper fraction or mixed number (do not reduce to lowest terms) .851

Parth (parthkohli):

.851 = 851/1000

OpenStudy (saifoo.khan):

\[\frac{0.851}{1} \to \frac{851}{1000}\]Simplify.